OpenStudy (ksaimouli):

A golfer hits the ball with a speed of 30 m/s. The ball touches the ground on the green 79.5 m away. At what approximate angle did the ball start? Neglect air drag

OpenStudy (ksaimouli):

@SithsAndGiggles

OpenStudy (anonymous):

\[\text{Range}=\frac{v^2\sin{2\theta}}{g}\] You're given range = 79.5 m v = 30 m/s g = acceleration due to gravity.

OpenStudy (ksaimouli):

A golfer hits the ball with a speed of 30 m/s. The ball touches the ground on the green 79.5 m away. Approximately, how long was the ball in the air? Neglect air drag

OpenStudy (ksaimouli):

i got that thx what about this

OpenStudy (anonymous):

Um, you just asked the same question again. Did you solve for theta?

OpenStudy (ksaimouli):

yes and this is not the same question i think

OpenStudy (ksaimouli):

i got 30

OpenStudy (anonymous):

Oh, sorry I only read the first half. Just a sec.

OpenStudy (anonymous):

I think the formula for air-time is something like \[t=\frac{2v\sin\theta}{g}\] I might be wrong; I haven't taken physics in 3 years.

OpenStudy (anonymous):

Use the formula: Range=( u^2 sin2x )/ g where x is the angle of projection.

OpenStudy (anonymous):

u- velocity, g- acc. due to gravity.